### v1

@@ Abstract
Fluid-structure interaction (FSI) simulation of biological systems presents significant computational challenges, particularly for applications involving large structural deformations and contact mechanics, such as heart valve dynamics. Traditional arbitrary Lagrangian-Eulerian methods encounter fundamental difficulties with such problems due to mesh distortion, motivating immersed techniques. This work presents a novel open-source immersed FSI framework that strategically couples two mature finite element libraries: MFEM, a GPU-ready and scalable library with state-of-the-art parallel performance developed at Lawrence Livermore National Laboratory, and FEBio, a nonlinear finite element solver with sophisticated solid mechanics capabilities designed for biomechanics applications developed at the University of Utah. This coupling creates a unique synergy wherein the fluid solver leverages MFEM’s distributed-memory parallelization and pathway to GPU acceleration, while the immersed solid exploits FEBio’s comprehensive suite of hyperelastic and viscoelastic constitutive models and advanced solid mechanics modeling targeted for biomechanics applications. FSI coupling is achieved using a fictitious domain/distributed Lagrange multiplier methodology with variational multiscale stabilization for enhanced accuracy on under-resolved grids expected with unfitted meshes used in immersed FSI. A fully implicit, monolithic scheme provides robust coupling for strongly coupled fluid-solid interactions characteristic of cardiovascular applications. The framework’s modular architecture facilitates straightforward extension to additional physics and element technologies. Several test problems are considered to demonstrate the capabilities of the proposed framework, including a three-dimensional semilunar heart valve simulation. This platform addresses a critical need for open-source immersed FSI software combining advanced biomechanics modeling with high-performance computing infrastructure.
@@
Fluid-structure interaction, Immersed finite element method, Immersed boundary, Fictitous domain, Biomechanics, Heart valve, MFEM, FEBio, Open-source software
@@ 1 Introduction
To address this need, we present a novel open-source immersed FSI framework coupling the modular finite element methods (MFEM) library [120], a scalable finite element library supporting massive parallelization via distributed memory computing and hardware accelerators (e.g. GPUs), with FEBio [121], a nonlinear finite element solver designed specifically for biomechanical applications. FEBio provides sophisticated solid mechanics capabilities including hyperelastic and viscoelastic constitutive models appropriate for soft tissues, robust contact algorithms essential for leaflet coaptation and tissue-device interaction, and specialized boundary conditions for cardiovascular analysis [121, 100, 101]. This immersed FSI framework overcomes the inherent limitations of ALE‑FSI methods for valve applications, enabling the study of pulsatile hemodynamics alongside the complex stress and strain fields in valve tissues necessary to understanding disease progression and mechanobiological tissue response. Although valve FSI modeling serves as the motivating application, the modular architecture of this framework accommodates diverse biomechanical and cardiovascular FSI applications. By providing an open-source platform with extensible workflows, this work aims to fill a critical gap in the computational biomechanics community, enabling integrated studies that address both fluid dynamic performance and tissue-level mechanical response across a broad spectrum of biomedical applications, including those with large solid deformations and motions.
@@ 2.1 Fluid
We discretize the above weak form using finite dimensional subspaces of the trial and test function spaces, and augment the equations with additional terms following the Variational Multiscale (VMS) method [85, 122]. We omit the VMS weak form for brevity and refer the reader to [65] for details. These stabilization terms and coefficients provide enhanced stability and accuracy on under-resolved grids (typical for unfitted FSI and advection-dominated flows), ensure optimal convergence rates with respect to mesh size and polynomial order, and allow for the use of computationally efficient equal-order interpolants for the velocity and pressure trial solution spaces [123]. Note the interpolant used for the Lagrange multiplier is discussed below in Section 2.3. Furthermore, we utilize a modification of the stabilization coefficients originally presented in [84] and leveraged in subsequent works [124, 78] as follows
@@ 3 Design of the MFEMiFSI Plugin
In addition to the fluid and solid contributions to the matrix system, the present methodology requires (parallel) computation of the FSI constraint matrix $\mathbf{C}^{\lambda}$. Recall from Equation 13, each row of the FSI constraint matrix corresponds to a fluid shape function evaluated at each solid node. Consequently, formation of this matrix requires a point search process and a non-standard assembly procedure. In the present work, we extend the existing parallel point search class in MFEM (FindPtsGSLIB), as well as leverage hypre’s IJ interface [143] and MFEM’s HypreParMatrix wrapper class to implement the formation of this matrix. FindPtsGSLIB provides an interface to efficiently communicate between MPI ranks where the point search process was started for a collection of local points and MPI ranks where these points were found. hypre’s IJ interface allows the user to interact with a distributed sparse matrix using row and column indices in the typical linear-algebra way. Pseudocode for our algorithm to compute $\mathbf{C}^{\lambda}$ is presented in Appendix A.
@@ 4 Numerical Test Problems
This section details the numerical benchmark and test problems that have been conducted to verify the implementation and demonstrate the capabilities of the immersed FSI framework. Note since FEBio is a three-dimensional (3D) finite element code, all two-dimensional (2D) problems are modeled in 3D with one element in the z direction and all solution components in the z direction set to zero. Unless stated otherwise, all example problems use first-order Lagrange polynomials as basis functions for the finite element approximations.
@@ 4.3 Oscillating flexible leaflet
We consider a set of three meshes for the fluid domain $32\times 192$, $64\times 384$, and $128\times 768$ with corresponding solid meshes $4\times 45$, $4\times 90$, and $6\times 180$ respectively. We simulate for 3 s using a time step size of $\Delta t=1\times 10^{-4}$ s, except for the finest mesh where we use $\Delta t=5\times 10^{-5}$ s for numerical stability. Several snapshots of the numerical solution over the final cycle are shown in Figure 9. Time histories of the leaflet tip displacement are shown in Figure 10. Overall, there is good agreement in the leaflet tip displacement between the present work and the reference solution, with some differences in the peak y-displacements that can be attributed to the use of different elements and different forms of solid constitutive models (uncoupled Neo-Hookean model in the present work). Furthermore, the authors from the reference solution utilized adaptive meshing techniques, making it difficult to reproduce the mesh resolution in the present work.
@@ 4.4 Free-falling sphere
Figure 13: Time history of the average sphere velocity in the gravitational direction under mesh refinement. Note the dashed black line denotes the analytical solution for a rigid free-falling sphere with the correction factor for the presence of rigid walls in our simulation.
@@ 4.5 Semilunar heart valve
In this section, we demonstrate the capability of the present framework to perform FSI simulation of three-dimensional heart valves immersed within a rigid blood vessel. Here, we simulate a tri-leaflet semilunar heart valve within an idealized aortic root. The present geometries, consisting of the leaflets and aortic root shown in Figure 14, are modification of the model provided from [153]. The fluid domain is $18.2$ cm in length and has a diameter of $2.8$ cm, as well as a three-lobed region near the valve to model the aortic root. The inlet, at the bottom of the geometry in Figure 14, is $4$ cm upstream of the root, while the outlet is $11.8$ cm downstream of the root. The fluid domain was meshed with $8,565,387$ tetrahedral elements, including boundary layer refinement near the vessel wall as seen in Figure 14, resulting in a total of $5,593,396$ DOFs for both the fluid velocity and pressure approximations. The leaflets have a height of $1.6$ cm and thickness of $0.015$ cm, and were meshed using a combination of hexahedral and pentahedral elements with four elements across the thickness resulting in $63,452$ elements and $215,970$ solid displacement DOFs.
@@
Figure 16 shows a series of snapshots of the leaflet motion and z-component of the fluid velocity over the simulated heartbeat. The flow field shows the development of a jet upon valve opening, followed by breakdown of the jet as the valve closes until forward flow essentially stops upon valve closure. Note the complex vortex dynamics downstream of the valve, which are expected flow features given the peak Reynolds number is approximately $5,400$ computed using the vessel diameter and peak velocity magnitude. We refer the reader to [95, 154] for discussions on turbulent flow features downstream of the valve. A top and isometric view of the leaflet displacement magnitude over the cardiac cycle is shown in Figure 17. Note the nonsymmetric deformations during valve closure, which are most likely attributed to the downstream nonsymmetric vortex dynamics. This asymmetric behavior is consistent with valve FSI simulations reported in [88, 84, 61].
@@
Figure 17: Snapshots of the leaflet displacement magnitude at several instances over the simulated heartbeat.
@@ 5 Discussion
The central contribution of this work is the coupling of two mature, open-source finite element libraries, MFEM and FEBio, each providing complementary capabilities essential for biological FSI applications. MFEM, developed at Lawrence Livermore National Laboratory, provides a scalable, GPU-ready infrastructure designed for exascale computing with support for distributed-memory parallelization as well as hardware accelerators (e.g. GPUs) enabling computing on both traditional and emerging architectures [120, 118]. FEBio, developed at the University of Utah, offers a comprehensive suite of biomechanics modeling features including hyperelastic and viscoelastic constitutive models appropriate for soft tissues, robust contact algorithms essential for leaflet coaptation, and specialized element formulations for nearly-incompressible materials [121]. This coupling creates a unique synergy: the fluid solver leverages MFEM’s parallel computing infrastructure while the immersed solid exploits FEBio’s wide range of constitutive modeling as well as advanced computational solid mechanics formulations. Unlike existing immersed FSI frameworks that require users to implement specialized biomechanics features, this approach provides these capabilities out of the box through FEBio’s established and validated codebase. While reduced order and simplified structural models decrease computational costs and can be effective for investigations focused on fluid performance, such approaches inadequately capture the tissue mechanics fundamental to understanding disease progression, calcific degeneration, mechanobiological response, and device-tissue interactions [41, 113, 114, 115, 116, 117]. The solid modeling capabilities provided via FEBio are essential for applications where tissue-level mechanics in combination with fluid dynamics drive clinical outcomes, including native valve repair optimization, bioprosthetic valve durability prediction, and transcatheter device evaluation [11, 47, 48].
@@
The immersed FSI method was implemented using FEBio’s plugin interface [141] and utilizes a modular design that facilitates straightforward extension to alternative solid formulations and additional modeling capabilities. The base class MFEMImmersedElasticDomain (Appendix B) defines virtual functions for computing element residual vectors and tangent matrices, enabling users to implement specialized element technologies, such as the mean dilatation method for locking-free nearly-incompressible behavior, without modifying core framework components. This design philosophy mirrors FEBio’s successful plugin framework, which has enabled community contributions spanning constitutive models, boundary conditions, and solver algorithms [141]. The modular structure also facilitates future integration of additional physics, including electromechanical coupling for cardiac simulations and reactive transport for thrombosis modeling. Furthermore, the framework inherits MFEM’s state-of-the-art parallel computing capabilities, enabling distributed-memory parallelization via MPI with a clear pathway to GPU acceleration [120]. MFEM has a large feature set, supporting a wide range of applications; discretizations types, including Galerkin, discontinuous Galerkin (DG), mixed methods, isogeometric analysis, hybridization, and discontinuous Petrov-Galerkin approaches; arbitrary high-order finite elements; the full de Rham complex of finite element spaces in 2D and 3D; a wide range of mesh types and support for mesh refinement and optimization; integration with external numerical libraries, such as PETSc [155] and hypre [156], for additional linear and nonlinear solvers; and massively parallel scalability, HPC efficiency, and GPU acceleration supporting a variety of architectures. We refer the reader to the following citations for more details [120, 118]. This infrastructure positions the framework to benefit from ongoing MFEM development efforts targeting exascale computing, including GPU-accelerated linear solvers and matrix-free operator evaluation as well as scalable automatic differentiation for complex applications [118, 157].
@@
Several limitations of the current framework should be acknowledged. Contact models should be incorporated into our solver to prevent intersections between immersed bodies rather than relying exclusively on FSI kinematics and the lubrication effect to prevent contact. FSI applications involving immersed thin structures, e.g. heart valve FSI analysis, could benefit from the inclusion of shell formulations to improve computational efficiency. Both contact and shell modeling are features readily available in FEBio and would need to be ported into the present framework, although this process is simplified by the modular design of the immersed solid solver described in Section 3. As an immersed (unfitted mesh) approach, the accuracy and efficiency of our computational FSI tool could be improved with adaptive mesh refinement (AMR) techniques to dynamically concentrate resolution near the fluid-solid interface. Accurate resolution of near-wall quantities is important for modeling a variety mechanobiological responses, such as vascular growth and remodeling [158, 159]. Additional performance gains could be realized by utilizing the existing GPU support in MFEM, which was not pursued in this work but is planned for future development. Lastly, the current implementation focuses on Newtonian fluids; extension to non-Newtonian models, while straightforward within the present framework, has not been completed.
Future development will focus on several key areas. Further parallelization and optimization of the plugin will be pursued leveraging the existing GPU support in MFEM, which will likely require further method, solver, and algorithmic development. Adaptive mesh refinement capabilities will be implemented to enable efficient resolution of flow features near the fluid-solid interface. FSI-specific contact algorithms will be developed to robustly handle leaflet coaptation during valve closure. Further integration between MFEM and FEBio will expand the range of biological FSI problems that can be addressed, including coupled electromechanical-FSI simulations and multiscale tissue modeling.

### v2

@@ Abstract
Fluid-structure interaction (FSI) simulation of biological systems presents significant computational challenges, particularly for applications involving large structural deformations and contact mechanics, such as heart valve dynamics. Traditional arbitrary Lagrangian-Eulerian methods encounter fundamental difficulties with such problems due to mesh distortion, motivating immersed techniques. This work presents a novel open-source immersed FSI framework that strategically couples two mature finite element libraries: MFEM, a GPU-ready and scalable library with state-of-the-art parallel performance developed at Lawrence Livermore National Laboratory, and FEBio, a nonlinear finite element solver with sophisticated solid mechanics capabilities designed for biomechanics applications developed at the University of Utah and Columbia University. This coupling creates a unique synergy wherein the fluid solver leverages MFEM’s distributed-memory parallelization and pathway to GPU acceleration, while the immersed solid exploits FEBio’s comprehensive suite of hyperelastic and viscoelastic constitutive models and advanced solid mechanics modeling targeted for biomechanics applications. FSI coupling is achieved using a fictitious domain/distributed Lagrange multiplier methodology with variational multiscale stabilization for enhanced accuracy on under-resolved grids expected with unfitted meshes used in immersed FSI. A fully implicit, monolithic scheme provides robust coupling for strongly coupled fluid-solid interactions characteristic of cardiovascular applications. The framework’s modular architecture facilitates straightforward extension to additional physics and element technologies. Several test problems are considered to demonstrate the capabilities of the proposed framework, including a three-dimensional semilunar heart valve simulation. This platform addresses a critical need for open-source immersed FSI software combining advanced biomechanics modeling with high-performance computing infrastructure.
@@
Fluid-structure interaction, Immersed method, Heart valve, MFEM, FEBio, Open-source software
@@ 1 Introduction
To address this need, we present a novel open-source immersed FSI framework coupling the modular finite element methods (MFEM) library [120], a scalable finite element library supporting massive parallelization via distributed memory computing and hardware accelerators (e.g. GPUs), with FEBio [121], a nonlinear finite element solver designed specifically for biomechanical applications. FEBio provides sophisticated solid mechanics capabilities including hyperelastic and viscoelastic constitutive models appropriate for soft tissues, robust contact algorithms essential for leaflet coaptation and tissue-device interaction, specialized boundary conditions for cardiovascular analysis, as well as capabilities for modeling transport, chemical reactions, growth, and remodeling [121, 100, 101, 122, 123, 124, 46]. This immersed FSI framework overcomes the inherent limitations of ALE‑FSI methods for valve applications, enabling the study of pulsatile hemodynamics alongside the complex stress and strain fields in valve tissues necessary to understanding disease progression and mechanobiological tissue response. Although valve FSI modeling serves as the motivating application, the modular architecture of this framework accommodates diverse biomechanical and cardiovascular FSI applications. By providing an open-source platform with extensible workflows, this work aims to fill a critical gap in the computational biomechanics community, enabling integrated studies that address both fluid dynamic performance and tissue-level mechanical response across a broad spectrum of biomedical applications, including those with large solid deformations and motions.
@@ 2.1 Fluid
We discretize the above weak form using finite dimensional subspaces of the trial and test function spaces, and augment the equations with additional terms following the Variational Multiscale (VMS) method [85, 125]. We omit the VMS weak form for brevity and refer the reader to [65] for details. These stabilization terms and coefficients provide enhanced stability and accuracy on under-resolved grids (typical for unfitted FSI and advection-dominated flows), ensure optimal convergence rates with respect to mesh size and polynomial order, and allow for the use of computationally efficient equal-order interpolants for the velocity and pressure trial solution spaces [126]. Furthermore, we utilize a modification of the stabilization coefficients originally presented in [84] and leveraged in subsequent works [127, 78] as follows
@@ 3 Design of the MFEMiFSI Plugin
In addition to the fluid and solid contributions to the matrix system, the present methodology requires (parallel) computation of the FSI constraint matrix $\mathbf{C}^{\lambda}$. Recall from Equation 13, each row of the FSI constraint matrix corresponds to a fluid shape function evaluated at each solid node. Consequently, formation of this matrix requires a point search process and a non-standard assembly procedure. In the present work, we extend the existing parallel point search class in MFEM FindPtsGSLIB [146], as well as leverage hypre’s IJ interface [147] and MFEM’s HypreParMatrix wrapper class to implement the formation of this matrix. FindPtsGSLIB provides an interface to efficiently communicate between MPI ranks where the point search process was started for a collection of local points and MPI ranks where these points were found. hypre’s IJ interface allows the user to interact with a distributed sparse matrix using row and column indices in the typical linear-algebra way. Pseudocode for our algorithm to compute $\mathbf{C}^{\lambda}$ is presented in Appendix A.
@@ 4 Numerical Test Problems
This section details the numerical benchmark and test problems that have been conducted to verify the implementation and demonstrate the capabilities of the immersed FSI framework. Note since FEBio is a three-dimensional (3D) finite element code, all two-dimensional (2D) problems are modeled in 3D with one element in the z direction and all solution components in the z direction set to zero. Unless stated otherwise, all example problems use first-order Lagrange polynomials as basis functions for the finite element approximations. Meshes for the following problems were constructed using FEBio Studio [121], Simvascular [97], or GMsh [148]. Numerical simulations reported in this work were performed on both the Respublica supercomputer at the Children’s Hospital of Philadelphia (CHOP) and the Anvil supercomputer at Purdue University.
@@ 4.3 Oscillating flexible leaflet
We consider a set of three meshes for the fluid domain $192\times 32$, $384\times 64$, and $768\times 128$ with corresponding solid meshes $4\times 45$, $4\times 90$, and $6\times 180$ respectively. We simulate for 3 s using a time step size of $\Delta t=1\times 10^{-4}$ s, except for the finest mesh where we use $\Delta t=5\times 10^{-5}$ s for numerical stability. Several snapshots of the numerical solution over the final cycle are shown in Figure 9. Time histories of the leaflet tip displacement are shown in Figure 10. Overall, there is good agreement in the leaflet tip displacement between the present work and the reference solution, with some differences in the peak y-displacements that can be attributed to the use of different elements and different forms of solid constitutive models (uncoupled Neo-Hookean model in the present work). Furthermore, the authors from the reference solution utilized adaptive meshing techniques, making it difficult to reproduce the mesh resolution in the present work.
@@ 4.4 Free-falling sphere
Figure 13: Time history of the average sphere velocity in the gravitational direction under mesh refinement. Further details on the sequence of meshes utilized in this example problem are provided in Table 1. Note the dashed black line denotes the analytical solution for a rigid free-falling sphere with the correction factor for the presence of rigid walls in our simulation.
@@ 4.5 Semilunar heart valve
In this section, we demonstrate the capability of the present framework to perform FSI simulation of three-dimensional heart valves immersed within a rigid blood vessel. Here, we simulate a tri-leaflet semilunar heart valve within an idealized aortic root. The present geometries, consisting of the leaflets and aortic root shown in Figure 14, are modifications of the model provided from [158]. The fluid domain is $18.2$ cm in length and has a diameter of $2.8$ cm, as well as a three-lobed region near the valve to model the aortic root. The inlet, at the bottom of the geometry in Figure 14, is $4$ cm upstream of the root, while the outlet is $11.8$ cm downstream of the root. The fluid domain was meshed with $8,565,387$ tetrahedral elements, including boundary layer refinement near the vessel wall as seen in Figure 14, resulting in a total of $5,593,396$ DOFs for both the fluid velocity and pressure approximations. The leaflets have a height of $1.6$ cm and thickness of $0.015$ cm, and were meshed using a combination of hexahedral and pentahedral elements with four elements across the thickness, resulting in $63,452$ elements and $215,970$ solid displacement DOFs.
@@
Figure 16 shows a series of snapshots of the leaflet motion and z-component of the fluid velocity over the simulated heartbeat. The flow field shows the development of a jet upon valve opening, followed by breakdown of the jet as the valve closes until forward flow essentially stops upon valve closure. Note the complex vortex dynamics downstream of the valve, which are expected flow features given the peak Reynolds number is approximately $5,400$ computed using the vessel diameter and peak velocity magnitude. We refer the reader to [95, 159] for discussions on turbulent flow features downstream of the valve. A top and isometric view of the leaflet 1st principal stress over the cardiac cycle is shown in Figure 17. Note the nonsymmetric deformations during valve closure, which are most likely attributed to the downstream nonsymmetric vortex dynamics. This asymmetric behavior is consistent with valve FSI simulations reported in [88, 84, 61].
@@
Figure 17: Snapshots of the leaflet 1st principal stress at several instances over the simulated heartbeat.
@@ 5 Discussion
The central contribution of this work is the coupling of two mature, open-source finite element libraries, MFEM and FEBio, each providing complementary capabilities essential for biological FSI applications. MFEM, developed at Lawrence Livermore National Laboratory, provides a scalable, GPU-ready infrastructure designed for exascale computing with support for distributed-memory parallelization as well as hardware accelerators (e.g. GPUs) enabling computing on both traditional and emerging architectures [120, 118]. FEBio, developed at the University of Utah and Columbia University, offers a comprehensive suite of biomechanics modeling features including hyperelastic and viscoelastic constitutive models appropriate for soft tissues, robust contact algorithms essential for leaflet coaptation, and specialized element formulations for nearly-incompressible materials [121]. This coupling creates a unique synergy: the fluid solver leverages MFEM’s parallel computing infrastructure while the immersed solid exploits FEBio’s wide range of constitutive modeling as well as advanced computational solid mechanics formulations. Unlike existing immersed FSI frameworks that require users to implement specialized biomechanics features, this approach provides these capabilities out of the box through FEBio’s established and validated codebase. While reduced order and simplified structural models decrease computational costs and can be effective for investigations focused on fluid performance, such approaches inadequately capture the tissue mechanics fundamental to understanding disease progression, calcific degeneration, mechanobiological response, and device-tissue interactions [41, 113, 114, 115, 116, 117]. The solid modeling capabilities provided via FEBio are essential for applications where tissue-level mechanics in combination with fluid dynamics drive clinical outcomes, including native valve repair optimization, bioprosthetic valve durability prediction, and transcatheter device evaluation [11, 47, 48].
@@
The immersed FSI method was implemented using FEBio’s plugin interface [144] and utilizes a modular design that facilitates straightforward extension to alternative solid formulations and additional modeling capabilities. The base class MFEMImmersedElasticDomain (Appendix B) defines virtual functions for computing element residual vectors and tangent matrices, enabling users to implement specialized element technologies, such as the mean dilatation method for locking-free nearly-incompressible behavior, without modifying core framework components. This design philosophy mirrors FEBio’s successful plugin framework, which has enabled community contributions spanning constitutive models, boundary conditions, and solver algorithms [144]. The modular structure also facilitates future integration of additional physics, including electromechanical coupling for cardiac simulations and reactive transport for thrombosis modeling. Furthermore, the framework inherits MFEM’s state-of-the-art parallel computing capabilities, enabling distributed-memory parallelization via MPI with a clear pathway to GPU acceleration [120]. MFEM has a large feature set, supporting a wide range of applications; discretizations types, including Galerkin, discontinuous Galerkin (DG), mixed methods, isogeometric analysis, hybridization, and discontinuous Petrov-Galerkin approaches; arbitrary high-order finite elements; the full de Rham complex of finite element spaces in 2D and 3D; a wide range of mesh types and support for mesh refinement and optimization; integration with external numerical libraries, such as PETSc [160] and hypre [161], for additional linear and nonlinear solvers; and massively parallel scalability, HPC efficiency, and GPU acceleration supporting a variety of architectures. We refer the reader to the following citations for more details [120, 118]. This infrastructure positions the framework to benefit from ongoing MFEM development efforts targeting exascale computing, including GPU-accelerated linear solvers and matrix-free operator evaluation as well as scalable automatic differentiation for complex applications [118, 162, 163].
@@
Several limitations of the current framework should be acknowledged. Contact models should be incorporated into our solver to prevent intersections between immersed bodies rather than relying exclusively on FSI kinematics and the lubrication effect to prevent contact. FSI applications involving immersed thin structures, e.g. heart valve FSI analysis, could benefit from the inclusion of shell formulations to improve computational efficiency. Both contact and shell modeling are features readily available in FEBio and would need to be ported into the present framework, although this process is simplified by the modular design of the immersed solid solver described in Section 3. As an immersed (unfitted mesh) approach, the accuracy and efficiency of our computational FSI tool could be improved with adaptive mesh refinement (AMR) techniques to dynamically concentrate resolution near the fluid-solid interface. Accurate resolution of near-wall quantities is important for modeling a variety mechanobiological responses, such as vascular growth and remodeling [164, 165]. Additional performance gains could be realized by utilizing the existing GPU support in MFEM, which was not pursued in this work but is planned for future development. Lastly, the current implementation focuses on Newtonian fluids; extension to non-Newtonian models, while straightforward within the present framework due to the existing support in FEBio, has not been completed.
Future development will focus on several key areas. Further parallelization and optimization of the plugin will be pursued leveraging the existing GPU support in MFEM, which will likely require further method, solver, and algorithmic development. Adaptive mesh refinement capabilities will be implemented to enable efficient resolution of flow features near the fluid-solid interface. Contact models, such as the potential based contact formulation existing in FEBio [166], will be incorporated into the framework and FSI-specific modifications will be investigated to robustly and accurately model seepage of fluid between the contacting bodies [167, 168, 169, 170, 171]. Further integration between MFEM and FEBio will expand the range of biological FSI problems that can be addressed, including coupled electromechanical-FSI simulations and multiscale tissue modeling.